OpenStudy (anonymous):

how do i graph this? sin t + 1/2 cos t = f(t)

OpenStudy (anonymous):

How about using a computer program?

OpenStudy (anonymous):

so i have to use a comp program for this?

OpenStudy (anonymous):

or you can write it as a single function of sine and graph that

OpenStudy (anonymous):

Or that, anyway, the computer will make it easier.

OpenStudy (anonymous):

how would I do that? make sin t = -1/2 cos t?

OpenStudy (anonymous):

that is for sure http://www.wolframalpha.com/input/?i=sin%28x%29%2B.5cos%28x%29

OpenStudy (anonymous):

guess what i really want is to understand how to do this. if i cant be done by hand then i want to at least understand the concept

OpenStudy (anonymous):

you can use \[a\sin(x)+b\cos(x)=\sqrt{a^2+b^2}\sin(x+\theta)\] where \[\tan(\theta)=\frac{b}{a}\]

OpenStudy (anonymous):

so the basic idea is that it will look just like sine, only with a different amplitude, and also a phase shift

OpenStudy (jamesj):

It can be done by hand and the identity sat73 wrote down is very useful in lots of contexts.

OpenStudy (anonymous):

thanks